# Non-Gaussian enhancement of precision in quantum batteries

### v1

@@ Appendix C Expectation values on non-Gaussian states
[…]
$\operatorname{Tr}\left[M\rho_{NG}\right]=\frac{1}{K}\operatorname{Tr}\left[a_{S_{m}}^{\dagger}\ldots\right.\left.a_{S_{1}}^{\dagger}a_{C_{n}}^{\#}\ldots a_{C_{1}}^{\#}a_{S_{1}}\ldots a_{S_{m}}\rho_{G}\right]$
where operators with subscript $S_{i}$ are those implementing the photon subtractions (analogous for additions by taking hermitian conjugates) on mode $i$ of the Gaussian state, and those with subscript $C_{j}$ comprise the expression of the observable $M$ whose expectation value we wish to calculate. From Wick’s theorem we know that the expectation value of a single product of ladder operators can be decomposed as the following sum:
$\operatorname{Tr}$ $\left[a_{S_{m}}^{\dagger}\ldots\right.\left.a_{S_{1}}^{\dagger}a_{C_{n}}^{\#}\ldots a_{C_{1}}^{\#}a_{S_{1}}\ldots a_{S_{m}}\rho_{G}\right]=$ (36)
$=$ $\sum_{\mathcal{P}}\prod_{\left\{\left(p_{1},\#\right),\left(p_{2},\#\right)\right\}\in\mathcal{P}}\operatorname{Tr}\left[a_{p_{1}}^{\#}a_{p_{2}}^{\#}\rho_{G}\right]$
[…]
@@
[…]
@@
Appendix D Example: kitten states
Cat states, defined in quantum optics as superpositions of coherent states with opposite phases, present significant challenges for experimental realization. However, it has been shown that these states can be approximated–hence the term kitten–by successively applying photon subtractions to a squeezed thermal state Ourjoumtsev et al. (2006); Wakui et al. (2007) (cat states would correspond to the limit of infinite subtractions). This approach not only simplifies their practical implementation but also yields a particular class of non-Gaussian distributions for which the statistical moments of the photon-number observable can be handled analytically.
@@ Appendix D Example: kitten states
$N_{cat}$ $=\frac{1}{K}\operatorname{Tr}[\ \underbrace{a^{\dagger}a}_{N}\ \underbrace{a\ldots a}_{m}\ \sigma\ \underbrace{a^{\dagger}\ldots a^{\dagger}}_{m}]\ =$ (39)
$=\quad\frac{1}{K}\operatorname{Tr}[\ \underbrace{a^{\dagger}\ldots a^{\dagger}}_{m+1}\ \underbrace{a\ldots a}_{m+1}\ \sigma]$
$N_{cat}^{2}=\frac{1}{K}\operatorname{Tr}[\ \underbrace{a^{\dagger}aa^{\dagger}a}_{N^{2}}\ \underbrace{a\ldots a}_{m}\ \sigma\ \underbrace{a^{\dagger}\ldots a^{\dagger}}_{m}]\ =$ (40)
$=\quad\frac{1}{K}\operatorname{Tr}[\ \underbrace{a^{\dagger}\ldots a^{\dagger}}_{m+1}\ aa^{\dagger}\ \underbrace{a\ldots a}_{m+1}\ \sigma]\ \stackrel{{\scriptstyle[a,a^{\dagger}]=1}}{{=}}$
@@
$K$ $=\operatorname{Tr}[\ \underbrace{a\ldots a}_{m}\ \sigma\ \underbrace{a^{\dagger}\ldots a^{\dagger}}_{m}]\ =$ (41)
$=\quad\operatorname{Tr}[\ \underbrace{a^{\dagger}\ldots a^{\dagger}}_{m}\ \underbrace{a\ldots a}_{m}\ \sigma]$
By examination of Eqs. 39-41, we notice in fact we only need to calculate
@@
Once the problem has been reduced to computing $f(r)$ using the method proposed in the previous section, we notice that the identification of perfect matchings of the ladder operators will differ depending on the parity of $r$. We will examine both alternatives separately.
Case $r$ even. In this case homogeneous matchings are possible (i.e matchings that pair all of the creation operators among themselves, and all of the annihilation operators among themselves).
  • The amount of different homogeneous matchings that can be done is
@@
  • If, on the contrary, we paired all of the creation operators with an annihilation one (all non-homogeneous pairs), we would have $r!$ possibilities (the first $a$ would choose among $r$ different $a^{\dagger}$’s, the second one among the remaining $r-1$, etc).
  • The mixed approach in this case would be to make some $a^{\dagger}a$ pairs (necessarily an even number $s=2j$ of them) and mix the rest of operators homogeneously. The number of options here takes the form:
@@
Taking into account all these possibilities and applying Eqs.36, 37, we obtain the expression:
@@
Case $r$ odd. In this case homogeneous matchings are not possible, and an odd number of operators of each kind have to be paired non-homogeneously. Reasoning analogously as in the even case, we are left with:
@@
Note that these expressions (Eqs.44,45) can be applied to compute the mean photon number of any Gaussian non-displaced state undergoing successive photon subtractions. However, unlike the case of kitten states (where the covariance matrix of the squeezed thermal state is diagonal), the equality $I_{1}=I_{2}$ will not necessarily hold in general.
Appendix E Minimum Ergotropy of non-Gaussian states
@@
where $m$ denotes their stellar rank, $U_{G}$ is a general Gaussian transformation, and $\rho^{0}_{\text{th}}$ is the passive thermal state of the system given some fixed fluctuation level parametrized by a temperature $T$. Figure 4 shows that in principle, higher $m$ (that is, higher non-gaussianity) results in an advantage in terms of our figure of merit $\Gamma$. However, every photon added to the system conveys an energy increase, which may result in a violation of the constraint $\delta N\leq\epsilon$, and thus in the impracticability of the resulting state for the application that was pursued. In what follows, we will derive the explicit expression for such increase.
@@
$\sigma^{0}_{th}=\bigg{(}\begin{array}[]{cc}\nu&0\\ 0&\nu\end{array}\bigg{)}\quad\quad\text{where}\quad\nu\equiv coth\big{(}\frac{\omega}{2T}\big{)}$
Since $\rho^{0}_{\text{th}}$ is Gaussian, its mean photon number, which we denote $N_{0}$, can be directly calculated through 23:
@@
Adding photons to this passive thermal state naturally conveys an increase in average energy. In particular, the least possible energy that one may reach after $m$ photon additions is that of the state $\rho_{\text{th}}^{(+m)}=a^{\dagger^{m}}\ \rho^{0}_{\text{th}}\ a^{m}$ (where we have dropped the indices since we are dealing with only one mode). That is, successively adding (analogous for subtracting) the $m$ photons to $\rho^{0}_{\text{th}}$, without previously performing any Gaussian transformations (which would increase energy even further). The analytical derivation of the mean photon number $\langle N\rangle=\langle a^{\dagger}a\rangle$ for this particular class of states $\rho_{\text{th}}^{(+m)}$ appears to be particularly simple. As before, this is equivalent to computing
@@
  • The central pair $a^{\dagger}a$ is fixed, and the other operators can be paired as $aa^{\dagger}$ in $m!$ different ways, all $m!$ cancelled with the denominator $K$. In this case we have just
@@
If some specific use-case has a very restrictive constraint $\epsilon$ on the maximum energetic input, the maximum allowed $m$ may be zero and so the correct charging unitary to use in our application can only be a Gaussian one. In Fig.6 one can see the optimal $\Gamma$ attainable with zero, one, two, and three photon-added states, as a function of the ergotropic constraint. Temperature has been set to $T=0.5K$ arbitrarily.
Figure 6: Optimal $\Gamma$ for Gaussian and non-gaussian photon-added single-mode states as a function of the constraint $\epsilon$, at fixed temperature $T=0.5K$
@@
Appendix F Effects of truncation
Despite the simplicity of the two particular cases outlined above, the exact calculations of expectation values on non-Gaussian states are, in general, computationally resource-demanding. They scale as $N!!$! with the number $N$ of ladder operators involved, counting both those included in the operator whose expectation value we are computing, and those applied to construct the non-Gaussian distribution. However, typical approaches to this exact calculation (i.e., discrete-variable approximations made by truncating the dimension of the Hilbert space) can lead to misleading and unfair conclusions, as we are about to show. We will proceed by proving that truncating the Hilbert space of a coherent state $|\alpha\rangle$, allows the violation of the Gaussian bound.
@@ Appendix F Effects of truncation
$|\alpha\rangle$ $=\sum_{n}\frac{\alpha^{n}}{n!}e^{-\frac{|\alpha|^{2}}{2}}|n\rangle\approx$ (52)
$\approx|\alpha\rangle_{t}$ $\equiv C\left(|0\rangle+\frac{\alpha}{\sqrt{1!}}|1\rangle+\frac{\alpha^{2}}{\sqrt{2!}}|2\rangle\right)$
@@
$C=\left(1+|\alpha|^{2}+\frac{|\alpha|^{4}}{2}\right)^{-1/2}.$ (53)
@@
which can be shown to be always greater than $|\alpha|$, which is the existing bound on $\Gamma$ for coherent states, as shown in Eq.12. In other words, by a mere approximation, we gain a non-classical resource in terms of charging precision. Obviously, by increasing the dimension of the Hilbert space enough, we should recover the correct result. However, this would make the calculations more tedious than the exact result we provide, employing the Wick theorem.

### v2

@@ Appendix C Appendix C: Expectation values on non-Gaussian states
where $\in\{\dagger$$\cdot\}$ denotes the type of ladder operator (creation or annihilation, respectively), and $K:=\operatorname{Tr}\left[\prod_{i=1}^{m}a_{i}^{\#}\rho_{G}(a_{i}^{\#})^{\dagger}\right]$ is a normalization factor that comes from the fact that ladder operators are non-unitary. By the cyclic property of the trace:
$\operatorname{Tr}\left[M\rho_{NG}\right]=\frac{1}{K}\operatorname{Tr}\left[a_{S_{m}}^{\dagger}\ldots\right.\left.a_{S_{1}}^{\dagger}a_{C_{n}}^{\#}\ldots a_{C_{1}}^{\#}a_{S_{1}}\ldots a_{S_{m}}\rho_{G}\right]$ (37)
where operators with subscript $S_{i}$ are those implementing the photon subtractions (analogous for additions by taking hermitian conjugates) on mode $i$ of the Gaussian state, and those with subscript $C_{j}$ comprise the expression of the observable $M$ whose expectation value we wish to calculate. From Wick’s theorem, we know that the expectation value of a single product of ladder operators can be decomposed as the following sum:
$\operatorname{Tr}\left[a_{S_{m}}^{\dagger}\ldots\right.\left.a_{S_{1}}^{\dagger}a_{C_{n}}^{\#}\ldots a_{C_{1}}^{\#}a_{S_{1}}\ldots a_{S_{m}}\rho_{G}\right]=\sum_{\mathcal{P}}\prod_{\left\{\left(p_{1},\#\right),\left(p_{2},\#\right)\right\}\in\mathcal{P}}\operatorname{Tr}\left[a_{p_{1}}^{\#}a_{p_{2}}^{\#}\rho_{G}\right]$ (38)
where $\mathcal{P}$ is the set of all possible perfect matchings of the $2m+n$ indices $\left(p_{k},\#\right)\in\left\{C_{1},\ldots,C_{n},S_{1},\ldots,S_{m}\right\}\times$ $\{\dagger,\cdot\}$ of the ladder operators, including loops (i.e. matchings of one of the operators with itself [loophaffnian]). All we are left with are expectation values of singlets or pairs of annihilation or creation operators on a Gaussian state, which can be computed by using the following identities:
@@
$I_{1}$ $=\operatorname{Tr}\left[a_{j}^{\dagger}a_{k}^{\dagger}\rho_{G}\right]=\frac{1}{4}\left[V_{jk}-V_{j+N,k+N}-i\left(V_{j,k+N}+V_{j+N,k}\right)\right]$
$I_{2}$ $=\operatorname{Tr}\left[a_{j}a_{k}\rho_{G}\right]=I_{1}^{*}$
$I_{3}$ $=\operatorname{Tr}\left[a_{j}^{\dagger}a_{k}\rho_{G}\right]=\frac{1}{4}\left[V_{jk}+V_{j+N,k+N}+i\left(V_{j,k+N}-V_{j+N,k}\right)-2\delta_{jk}\right]$
$I_{4}$ $=\operatorname{Tr}\left[a_{j}a_{k}^{\dagger}\rho_{G}\right]=\delta_{jk}+I_{3}^{*}$
@@
Appendix D Appendix D: Kitten states example
Cat states, defined in quantum optics as superpositions of coherent states with opposite phases, present significant challenges for experimental realization. However, it has been shown that these states can be approximated–hence the term kitten–by successively applying photon subtractions to a squeezed thermal state [kitten_states, Wakui2007] (cat states would correspond to the limit of infinite subtractions). This approach not only simplifies their practical implementation but also yields a particular class of non-Gaussian distributions for which the statistical moments of the photon-number observable can be handled analytically.
@@ Appendix D Appendix D: Kitten states example
$N_{cat}=\frac{1}{K}\operatorname{Tr}[\ \underbrace{a^{\dagger}a}_{N}\ \underbrace{a\ldots a}_{m}\ \sigma\ \underbrace{a^{\dagger}\ldots a^{\dagger}}_{m}]\ =\quad\frac{1}{K}\operatorname{Tr}[\ \underbrace{a^{\dagger}\ldots a^{\dagger}}_{m+1}\ \underbrace{a\ldots a}_{m+1}\ \sigma]$ (41)
$N_{cat}^{2}$ $=\frac{1}{K}\operatorname{Tr}[\ \underbrace{a^{\dagger}aa^{\dagger}a}_{N^{2}}\ \underbrace{a\ldots a}_{m}\ \sigma\ \underbrace{a^{\dagger}\ldots a^{\dagger}}_{m}]\ =\quad\frac{1}{K}\operatorname{Tr}[\ \underbrace{a^{\dagger}\ldots a^{\dagger}}_{m+1}\ aa^{\dagger}\ \underbrace{a\ldots a}_{m+1}\ \sigma]\ \stackrel{{\scriptstyle[a,a^{\dagger}]=1}}{{=}}$ (42)
@@
$K=\operatorname{Tr}[\ \underbrace{a\ldots a}_{m}\ \sigma\ \underbrace{a^{\dagger}\ldots a^{\dagger}}_{m}]\ =\quad\operatorname{Tr}[\ \underbrace{a^{\dagger}\ldots a^{\dagger}}_{m}\ \underbrace{a\ldots a}_{m}\ \sigma]$ (43)
By examination of Eqs. 41-43, we notice in fact we only need to calculate
@@
Once the problem has been reduced to computing $f(r)$ using the method proposed in the previous section, we observe that the identification of perfect matchings of the ladder operators depends on the parity of $r$. We will examine both alternatives separately.
Case $r$ even. In this case, homogeneous matchings are possible (i.e, matchings that pair all of the creation operators among themselves, and all of the annihilation operators among themselves).
  • The number of different homogeneous matchings that can be done is
@@
  • If, on the contrary, we paired all of the creation operators with an annihilation one (all non-homogeneous pairs), we would have $r!$ possibilities (the first $a$ would choose among $r$ different $a^{\dagger}$’s, the second one among the remaining r1r-1, etc).
  • The mixed approach in this case would be to make some $a^{\dagger}a$ pairs (necessarily an even number $s=2j$ of them) and mix the rest of the operators homogeneously. The number of options here takes the form:
@@
Taking into account all these possibilities and applying Eqs.38, 39, we obtain the expression:
@@
Case $r$ odd. In this case, homogeneous matchings are not possible, and an odd number of operators of each kind have to be paired non-homogeneously. Reasoning analogously as in the even case, we are left with:
@@
Note that these expressions (Eqs.46,47) can be applied to compute the mean photon number of any Gaussian non-displaced state undergoing successive photon subtractions. However, unlike the case of kitten states (where the covariance matrix of the squeezed thermal state is diagonal), the equality $I_{1}=I_{2}$ will not necessarily hold in general.
Appendix E Appendix E: Minimum Ergotropy of non-Gaussian states
@@
where $m$ denotes their stellar rank, $U_{G}$ is a general Gaussian transformation, and $\rho^{0}_{\text{th}}$ is the passive thermal state of the system given some fixed fluctuation level parametrized by a temperature $T$. Figure 4 in the main text shows that, in principle, higher $m$ (that is, higher non-gaussianity) results in an advantage in terms of our figure of merit $\Gamma$. However, every photon added to the system conveys an energy increase, which may result in a violation of the constraint $\delta N\leq\epsilon$, and thus in the impracticability of the resulting state for the application that was pursued. In what follows, we will derive the explicit expression for such an increase.
@@
[…]
Since $\rho^{0}_{\text{th}}$ is Gaussian, its mean photon number, which we denote $N_{0}$, can be directly calculated through 25:
@@
Adding photons to this passive thermal state naturally increases the average energy. In particular, the least possible energy that one may reach after $m$ photon additions is that of the state $\rho_{\text{th}}^{(+m)}=a^{\dagger^{m}}\ \rho^{0}_{\text{th}}\ a^{m}$ (where we have dropped the indices since we are dealing with only one mode). That is, successively adding (analogous for subtracting) the $m$ photons to $\rho^{0}_{\text{th}}$, without previously performing any Gaussian transformations (which would increase energy even further). The analytical derivation of the mean photon number $\langle N\rangle=\langle a^{\dagger}a\rangle$ for this particular class of states $\rho_{\text{th}}^{(+m)}$ appears to be particularly simple. As before, this is equivalent to computing
@@
  • The central pair $a^{\dagger}a$ is fixed, and the other operators can be paired as $aa^{\dagger}$ in $m!$ different ways, all $m!$ cancelled with the denominator $K$. In this case, we have just
@@
If some specific use-case has a very restrictive constraint $\epsilon$ on the maximum energetic input, the maximum allowed $m$ may be zero, and so the correct charging unitary to use in our application can only be a Gaussian one. In Fig.6, one can see the optimal $\Gamma$ attainable with zero, one, two, and three photon-added states, as a function of the ergotropic constraint. The temperature has been set to $T=0.5K$ arbitrarily.
Figure 6: Optimal $\Gamma$ for Gaussian and non-Gaussian photon-added single-mode states as a function of the constraint $\epsilon$, at fixed temperature $T=0.5K$.
@@
Appendix F Appendix F: Effects of truncation
Despite the simplicity of the two particular cases outlined above, the exact calculations of expectation values on non-Gaussian states are, in general, computationally resource-demanding. They scale as $n!!$ with the number $n$ of ladder operators involved, counting both those included in the operator whose expectation value we are computing, and those applied to construct the non-Gaussian distribution. However, typical approaches to this exact calculation (i.e., discrete-variable approximations made by truncating the dimension of the Hilbert space) can lead to misleading and unfair conclusions, as we are about to show. We will proceed by proving that truncating the Hilbert space of a coherent state $|\alpha\rangle$, allows the violation of the Gaussian bound.
@@ Appendix F Appendix F: Effects of truncation
$|\alpha\rangle=\sum_{n}\frac{\alpha^{n}}{n!}e^{-\frac{|\alpha|^{2}}{2}}|n\rangle\approx\underbrace{C\left(|0\rangle+\frac{\alpha}{\sqrt{1!}}|1\rangle+\frac{\alpha^{2}}{\sqrt{2!}}|2\rangle\right)}_{\equiv\ |\alpha\rangle_{t}}$ (54)
@@
$C=\left(1+|\alpha|^{2}+\frac{|\alpha|^{4}}{2}\right)^{-1/2}$ (55)
@@
which can be shown to be always greater than $|\alpha|$, which is the existing bound on $\Gamma$ for coherent states, as shown in Eq.12 in the main text. In other words, by a mere approximation, we gain a non-classical resource in terms of charging precision. Obviously, by increasing the dimension of the Hilbert space enough, we should recover the correct result. However, this would make the calculations more tedious than the exact result we provide, employing the Wick theorem.
Appendix G Appendix G: Multimode case
Lemma 1 establishes the independency of the signal-to-noise ratio $\Gamma$ under Gaussian global (entangling) passive transformations in multimode Gaussian states. Below we provide a proof of this result:
Proof. Consider a Gaussian state of $m$ modes $\rho$. Our figure-of-merit $\Gamma(\rho)$ is defined as the quotient between the difference in photon number in $\rho$ as compared to its thermal passive state $\rho_{0}$, and the photon number uncertainty $\Delta N(\rho)$:
$\Gamma(\rho)=\frac{N(\rho)-N(\rho_{0})}{\Delta N(\rho)}.$ (57)
Let $O$ be some generic passive-optics Gaussian transformation. Then $O\in O(2N)\cap Sp_{2n,\mathds{R}}$, i.e. $O$ is an orthogonal and symplectic matrix. By definition of passitivity, the application of $O$ cannot change the energy (nor the photon number) of state $\rho$, and thus the numerator is trivially not affected. As for the denominator:
$\Delta N(\rho)=\sqrt{\langle N^{2}\rangle_{\rho}-N(\rho)^{2}},$ (58)
where $N(\rho)^{2}$ is invariant since, as argued above, $N(\rho)$ is. The only term that could be possibly affected by the application of $O$ is then $\langle N^{2}\rangle_{\rho}$. Denoting the vector of quadratures $\bm{R}=(x_{1},\ldots x_{m},p_{1},\ldots p_{m})^{T}$, we expand $\langle N^{2}\rangle_{\rho}$ in terms of $\rho$’s covariance matrix $\sigma$ and displacement vector $\bm{\alpha}$:
$\langle N^{2}\rangle=\langle\bigg(\sum_{i=1}^{m}a^{\dagger}_{i}a_{i}\bigg)^{2}\rangle=\langle\bigg(\sum_{i=1}^{m}\frac{1}{2}(x_{i}^{2}+p_{i}^{2}-1)\bigg)^{2}\rangle=\langle\bigg(\frac{1}{2}(\bm{R}^{T}\bm{R}-m)\bigg)^{2}\rangle=\frac{1}{4}(\langle(\bm{R}^{T}\bm{R})^{2}\rangle-2m\langle\bm{R}^{T}\bm{R}\rangle+m^{2})$ (59)
By definition, $\bm{\alpha=\langle R\rangle}$, and $\sigma_{ij}=\frac{1}{2}(\{r_{i},r_{j}\}),$ where $\bm{r}\equiv\bm{R}-\bm{\alpha}$, with $\langle\bm{r}\rangle=0$. Hence:
[…]
While the expectation value of its square yields:
$\langle(\bm{R}^{T}\bm{R})^{2}\rangle=\langle(\bm{\alpha}^{T}\bm{\alpha})^{2}\rangle+2\langle\bm{\alpha}^{T}\bm{r}\bm{r}^{T}\bm{\alpha}\rangle+\langle(\bm{\alpha}^{T}\bm{r})^{2}\rangle+\langle(\bm{r}^{T}\bm{\alpha})^{2}\rangle+\langle(\bm{r}^{T}\bm{r})^{2}\rangle+4\langle(\bm{\alpha}^{T}\bm{\alpha})(\bm{\alpha}^{T}\bm{r})\rangle+2\langle(\bm{\alpha}^{T}\bm{\alpha})(\bm{r}^{T}\bm{r})\rangle+4\langle(\bm{\alpha}^{T}\bm{r})(\bm{r}^{T}\bm{r})\rangle$ (61)
Terms in the previous expression with an odd number of $\bm{r}$’s cancel, and we are left with:
[…]
Employing Wick’s theorem, the term $\langle(\bm{r}^{T}\bm{r})^{2}\rangle$ can be expressed as
$\langle(\bm{r}^{T}\bm{r})^{2}\rangle=2\text{Tr}(\sigma^{2})+(\text{Tr}(\sigma))^{2}-m$ (63)
By inspection of Eqs.60-63, it suffices to prove that $\bm{\alpha}||$, $\bm{\alpha}^{T}\sigma\bm{\alpha}$, $\text{Tr}(\sigma)$, and $\text{Tr}(\sigma^{2})$ are invariant under the application of passive optics transformations:
[…]
Since all of its components are invariant under $O$, $\Gamma(\rho)$ also remains unchanged. ∎
In Fig.7 we show the value of $\Gamma$ for the optimal two-mode Gaussian state (dashed black), and one-photon added state, both in the separable case (light blue, yielding the worst ratio) and the maximally entangled one (dark blue). The shaded area represents the gain given by the application of a $\theta=\pi/4$ beam splitter. Note that, at low temperatures, entanglement plays no role, since, as mentioned above, in the pure-state regime, Fock states are always optimal in terms of photon-number uncertainty.